Shift Summary for Owl shift, Thu. 1/21/2000


Subject: Shift Summary for Owl shift, Thu. 1/21/2000
From: Vivian O'Dell (odell@fnal.gov)
Date: Fri Feb 04 2000 - 15:34:52 CST


              Shift Summary for Owl shift, Thu. 1/21/2000

On Shift: PMcB, SG and many others

Hours of beam available at all intensities: 7.0
Hours of beam available at nominal* intensity: 7.0
Hours of KAMI test data: 4.5

 
Accounting of lost time & intensity:

        Access to change configuration

========================================================================
Runs ended this shift:
  15647 -- KAMI_KE3
  15648 -- KAMI_MUON
  15650 -- KAMI_KE3 4e12 20"Be (no disk for east beam data)
  15651 -- KAMI_KE3 4e12 0" Be
  15652 -- KAMI_KE3 5e11 20"Be
  15653 -- KAMI_HADRON 5e11 20"Be
  15654 -- KAMI_HADRON 5e11 0"Be
  
  
========================================================================
Tapes brought to FCC this shift: None.
========================================================================
Accesses during shift:

  Access to change configuration after AGEL runs
========================================================================
Overview of shift:

  Completed AGEL ke3 runs
  Completed KAMI_KE3 and a couple of short KAMI_HADRON runs

========================================================================
Description of Shift
 0:00 call MCR to change request to 4e12

 0:40 AGEL data set finishing; page BobH

 1:25 Call for access to change configuration
                TRD's are now OUT and OFF
                BAG and UCAL are now both in EAST Hole
                Setup single quartz bar + phototube behind BAG

 2:30 DJ and Wah removed all poly from neutral beam in NM2

 4:00 Muon run 15648 for DC alignment

 4:20 Run 15650 started (KAMI_KE3 4e12, 20" Be)

 5:00 Start run 15651 (KAMI_KE3 4e12 0" Be)
        See evidence that UCAL sags @ 4 MHz

 5:45 Run 15652 (KAMI_KE3 5e11 20"Be)

 6:50 Run 15653 KAMI_HADRON uses UCAL self trigger to study neutrons in BAG
and BAH
                  a few spills (5e11 20" Be)
 7:00 Run 15654 KAMI_HADRON UCAL self trigger (5e11 0" Be) a few spills

 7:00 End KAMI_KE3 runs
        Hand over to new shift crew

    



This archive was generated by hypermail 2b27 : Fri Feb 04 2000 - 15:34:55 CST